OpenStudy (britbrat1997):

Find the equation (in terms of x) of the line through the points (-4,3) and (2,1)

OpenStudy (fibonaccichick666):

So, what's the slope formula?

OpenStudy (britbrat1997):

y=mx+b is that the correct one or is it asking you y= y2-y1/x2-x1?

OpenStudy (fibonaccichick666):

the second one

OpenStudy (fibonaccichick666):

very good, ok, so do you know how to calculate the slope? (find m)

OpenStudy (britbrat1997):

i just use the formula( plug i n the points) and my answer is m, is that correct? if not can you tell me the right way?

OpenStudy (fibonaccichick666):

Oh and it's m not y. \[m=\frac{ y_2-y_1}{x_2-x_1}\]

OpenStudy (fibonaccichick666):

And yep, so what is your y1?

OpenStudy (britbrat1997):

give me one moment and ill work it out

OpenStudy (fibonaccichick666):

or just show the work and I can check it

OpenStudy (fibonaccichick666):

I need you to show your steps so I can make sure you did it right, ok?

OpenStudy (britbrat1997):

okay

OpenStudy (britbrat1997):

[m=1-3\div2-(-4)\]

OpenStudy (britbrat1997):

m=1-3/2-(-4)

OpenStudy (britbrat1997):

-2/6=-0.333 is this correct?

OpenStudy (fibonaccichick666):

very good, now remember you can't just plug that into a calculator like that

OpenStudy (fibonaccichick666):

yep yep

OpenStudy (fibonaccichick666):

so, let's just reduce that fraction

OpenStudy (britbrat1997):

i did 1-3 first then2--4 then i divided the outcomes or -2/6

OpenStudy (fibonaccichick666):

what can you divide both 2 and 6 by evenly?

OpenStudy (fibonaccichick666):

That's perfect

OpenStudy (britbrat1997):

1/3?

OpenStudy (fibonaccichick666):

ok good, so it still stays negative though

OpenStudy (britbrat1997):

okay so the final answer would be -1/3 ?

OpenStudy (fibonaccichick666):

that's the m (slope)

OpenStudy (fibonaccichick666):

we still have to come up with an equation

OpenStudy (britbrat1997):

since its asking for the equation, how should i go about that?

OpenStudy (britbrat1997):

y=1/3x+b do i need to come up with the b?

OpenStudy (fibonaccichick666):

So two ways, 1: you were given something called point slope form 2: we do the quick derivation

OpenStudy (fibonaccichick666):

3: you graph it and hope your y intercept is a nice number

OpenStudy (britbrat1997):

um can you help me with whatever one you think is best? im not sure how to do any tbh. what do i graph?

OpenStudy (fibonaccichick666):

ok, so let's do the very quick derivation, just so you get to see it

OpenStudy (fibonaccichick666):

So before, you told me that we find slope by \[m=\frac{y_2-y_1}{x_2-x_1}\] So if we wanted to get the ys by themselves, we could multiply both sides by the bottom of the fraction and 'undo' the division. \[m(x_2-x_1)=\frac{y_2-y_1}{x_2-x_1}*(x_2-x_1)\] So the right hand side cancels and we get: \[m(x_2-x_1)=y_2-y_1\] Since we prefer the y's on the left hand side we flip it, \[y_2-y_1=m(x_2-x_1)\] And we drop the sub twos \[y-y_1=m(x-x_1)\] We now have point slope form :)

OpenStudy (fibonaccichick666):

So for point slope form, if you have a point you put it in place of x1 and y1. Then you found m already, so you are all set to simplify